For the reaction system: 2NO(g) + O2(g) ? 2NO2(g) volume is suddenly reduce to half its value by increas- ing the pressure on it. If the reaction is of first order with respect to O2 and second order with respect to NO, the rate of reaction will

Grade:12

2 Answers

Naveen Kumar
9 years ago
2NO(g) + O2(g) …............> 2NO2(g)
ratewe is given as: r=k[NO]x * [O2]y
where rate with respect to No=x(=2,given)
and with O2, it is=y(=1,given)
Rate of reaction is given as overall sum of the order with ndividual components,
so orderof reaction=x+y=1+2=3
so r=k[NO]2* [O2]
After decreasing the volume, the concentration of the species will increase in same proportion
r’=k{2*[NO]}2 *2*[O2]=r*8
so it will be inceased by 8 times of the previous rate.
ankit singh
3 years ago
2NO left parenthesis straight g right parenthesis space plus space straight O subscript 2 left ... if the volume of the reaction vessel is diminished to one third of its original volume? Will there be any change in the order of the reaction with the reduced volume ? The below reaction take place in one step, thus rate of reaction is